OpenStudy (loser66):

I use wolframalpha to check my solution, however, I cannot open step by step solution. Anyone guide me, please

OpenStudy (anonymous):

~Do you have a account on wolframlpha?~

OpenStudy (loser66):

sure.

OpenStudy (anonymous):

Have you used up your 3 for the day?

OpenStudy (loser66):

what do you mean? I paid monthly

OpenStudy (anonymous):

Hmm, nevermind. Well, maybe you should restart your computer and see if it works? Maybe your computer Is being a little laggy and not opening it up.

OpenStudy (loser66):

ok, thanks.

OpenStudy (anonymous):

Welcome.

OpenStudy (amistre64):

step by step for what?

OpenStudy (loser66):

I typed my problem on the question bar, and it just gives me the final answer. I click "step by step solution" it dropped out the form but nothing there

OpenStudy (anonymous):

@Loser66, have you learned about the Laplace transform?

OpenStudy (loser66):

yes, I was taught this morning

OpenStudy (loser66):

but for this problem, I was asked to solve by U.C and variation parameter, both ways are not hard but take time. I want to check my answer

OpenStudy (anonymous):

Both methods, or just one of them? U.C. is simpler, I think.

OpenStudy (loser66):

Both

OpenStudy (loser66):

Ha, variation parameter is nightmare, hehehe. but still has an outlet.

OpenStudy (loser66):

I send message to wolframalpha to get help from the site. do you know how to fix my problem from the site.?

OpenStudy (anonymous):

I've opened the link, but I don't get a solution nor do I have the option to see the step-by-step solution. I could help you check your answer if you'd like.

OpenStudy (loser66):

let me scan my solution , it is more than 2 pages, hehe

OpenStudy (loser66):

nope, characteristic is r^3 -1=0

OpenStudy (anonymous):

Ah, right. Disregard that comment.

OpenStudy (anonymous):

In the first scan, when you substituted \(y_{p_1}\) and \(y'''_{p_1}\), you should have \[3Ae^{t}+Ate^t-Ate^t=\frac{1}{2}e^t\\ 3A=\frac{1}{2}\\ A=\frac{1}{6}\] (it looks like you had it right at one time)

OpenStudy (loser66):

yes, thanks a lot. you are right

OpenStudy (loser66):

yes, I think you pointed out my only one mistake which leads to wrong answer compare to wolfram.

OpenStudy (loser66):

I have many problems else. when I finish, I will post here. If you have time, please stop by and check it. It's not urgent. Whenever you have time, check one by one, please. I need it.

OpenStudy (anonymous):

Sure thing

OpenStudy (loser66):

Thanks a looooot.

OpenStudy (anonymous):

You're welcome!